If a right circular cylindrical vessel is one-quarter full, and you added seven liters of liquid that made the vessel three-fifths full, what is the total capacity of the vessel in liters?
Answer by rfer(16322)

3/5-1/4=
12/20-5/20=7/20
7/20=7 liters
7/(7/20)=x
7*20/7=20 liters